OpenStudy (anonymous):

"given f(x)=x^2 what is the domain of g(x)=f(x+2)-5"

OpenStudy (rsadhvika):

Any kind of transformation on a parabola will not change its domain.

OpenStudy (anonymous):

this is a composition isn't it?

OpenStudy (rsadhvika):

Having said that, g(x) domain would be same as f(x)

OpenStudy (rsadhvika):

It is a composition, you can look at it as transformation also.

OpenStudy (rsadhvika):

Oops ! I was wrong

OpenStudy (rsadhvika):

We need to work this as composition only

OpenStudy (anonymous):

so it wouldn't just be the domain of: g(f(x))=(x^2 +2)-5

OpenStudy (anonymous):

is that the proper composition? and then you find the domain that way?

OpenStudy (rsadhvika):

There is a trick, which simplifies finding the domain

OpenStudy (rsadhvika):

g(x)=f(x+2)-5

OpenStudy (rsadhvika):

what values we feeding in to g(x) ? Those all are output given by f only

OpenStudy (anonymous):

so x is in the domain of real numbers with no restrictions? like x^2?

OpenStudy (rsadhvika):

g(x) never gets a value less than 2 to its input.

OpenStudy (rsadhvika):

Domain of g(x) : x > = 2

OpenStudy (anonymous):

why???

OpenStudy (anonymous):

um this is kind of confusing. would you mind just telling me if I composed the function properly?

OpenStudy (anonymous):

g(f(x))=(x^2 +2)-5

OpenStudy (rsadhvika):

hold on the composition for a bit. Domain of f(x) : all reals Range of f(x) : x > 0 Domain of g(x) is the Range of f(x)

OpenStudy (rsadhvika):

On second thoughts, Domain of g(x) is : x >= 0, which is the range of f(x)

OpenStudy (rsadhvika):

Still its confusing ?

OpenStudy (cwrw238):

g(x) = (x + 2)^2 - 5 = x^2 + 4x - 1 the domain of f(x) is all real x the domain of g(x) is the same as f(x)

OpenStudy (rsadhvika):

thats wrong cwrw, we need to see f(x) and g(x) as composition. your working makes g(x) isolated

OpenStudy (cwrw238):

the range of g(x) is [-5, + infinity)

OpenStudy (rsadhvika):

g(x) never gets a negative value at its input because, all values it gets in input are fed from the output of f

OpenStudy (cwrw238):

yes i see your point - but the question is somewhat ambiguous

OpenStudy (cwrw238):

i'm confused also lol

OpenStudy (rsadhvika):

^agree it seems ambiguous first read

OpenStudy (rsadhvika):

I twisted the answer so many times before seeing it a bit clearly.. seems Cutie ran away lol

OpenStudy (cwrw238):

yea

OpenStudy (cwrw238):

can g(x) be a composition? if so whats the function h in g(x) = h(f(x) ?

OpenStudy (rsadhvika):

i would see it like this proly : h is a simply composition g is a mix of composition and transformation

OpenStudy (rsadhvika):

*simple

OpenStudy (anonymous):

sorry, i was really just confused about the wording too. does the way the f was placed mean the question is really what is the domain of g(fx) or something else?

OpenStudy (rsadhvika):

we can see it like this :- let h(x) = x+2 g(x) = f(h(x)) - 5

OpenStudy (rsadhvika):

we still need to find domain of g(x) only

OpenStudy (rsadhvika):

if you think f(h(x)) as variable t g(x) = t-5

OpenStudy (rsadhvika):

t can take only values : t >= 0

OpenStudy (rsadhvika):

so domain is [0, +inf)

OpenStudy (cwrw238):

i've just input the problem into the wolfram alpha math software program. it ignored the f\9x) = x^2 bit and came back with the domain of g(x) being all real numbers i think theres something wrong with the question

OpenStudy (rsadhvika):

interesting, can you give the link

OpenStudy (john_es):

The domain is all real numbers, g(x) is defined for all inputs x in R.

OpenStudy (rsadhvika):

question clearly says g is a composition of f

OpenStudy (cwrw238):

hhmmm - i dont think so

OpenStudy (john_es):

Even if it is a composition, the domain is still all R. However if g(x)=Log(x) it would be another story,

OpenStudy (rsadhvika):

would like to see wat @UnkleRhaukus thinks

OpenStudy (dumbcow):

by definition \[f(x+2) = (x+2)^{2}\] thus \[g(x) = (x+2)^{2} -5\] so domain of g is all real numbers @rsadhvika had it right at first then a lot of over analyzing i guess

OpenStudy (anonymous):

f(x) is a polynomial function whose domain is all reals: